### v2

# Emergent broadband polarization entanglement from electronic and phononic Stokes–anti-Stokes indistinguishability

Abstract
Recently [PRA 108, L051501 (2023)], it has been shown that in a centrosymmetric cubic system, two-photons from a broadband intense laser field can be converted into a pair of Stokes and anti-Stokes (SaS) entangled photons. While the previous work was based on symmetry arguments, here we present a fully quantum theory for the SaS scattering that properly explains, quantitatively describes, and provides a means to predict its spectral and polarization properties (for diamond). We also explore the possibilities offered by such system, designing an entanglement map based on changes in the light-matter system. In particular, we show how the broadband polarization entanglement, that emerges from the interference between electronic and phononic degrees of freedom in the SaS scattering, depends on parameters such as Stokes-anti-Stokes Raman shift, scattering geometry and laser bandwidth, opening the avenue of exploration of such phenomenon in information processing.
Photons produced in Raman scattering have their frequency shifted in relation to the incoming light ($\omega_{L}$), being called Stokes if their frequency is red-shifted ($\omega_{S}$) and anti-Stokes if they are blue-shifted ($\omega_{aS}$). The two of them can be produced as part of a higher order process, which results in their correlation ($\hbar\omega_{L}+\hbar\omega_{L}^{\prime}=\hbar\omega_{S}+\hbar\omega_{aS}$), with the coupled quanta being known as the SaS photon pair [1, 2, 3]. The operation of converting incoming photon(s) into different outgoing photon(s) in a medium is dictated by its susceptibility $\chi$ [4, 5], and the $\chi^{(3)}$ enables two photons from a broadband intense laser field to interact with the medium to produce the correlated SaS photon pair. In centrosymmetric materials, the second-order susceptibility is null and the non-linear response is usually dominated by the third-order tensor [6]. While the non-classical correlations between SaS photons have been widely established [7, 8, 9, 10, 11], the presence of polarization entanglement in such process has been demonstrated only very recently, in diamond [12].
This recent result indicates that the SaS polarization entanglement comes from the indistinguishability between electronic (e-SaS) and phononic (p-SaS) mediated processes inside the centrosymmetric material. It should, therefore, be possible to tailor the SaS state by playing with state preparation parameters, such as the scattering geometry, the SaS Raman shift and the laser bandwidth [13, 14, 15]. In fact, the interplay of the SaS scattering properties should generate a map of different degrees of polarization entanglement depending on the balance between the electronic and phononic degrees of freedom. Properly demonstrating, quantifying and exploring these possibilities is the goal of this paper. To achieve this, this work presents and experimentally certifies a fully quantum theory that describes the broadband scattered state in the SaS process, enabling a whole new level of the engineering of SaS photon pair generation.
The correlated SaS scattered intensity can be obtained by the electric polarization in the material [4, 5]
$P_{i}(\omega_{aS})\propto\chi^{(3)}_{ijkl}(-\omega_{aS},\omega_{L},\omega_{L}^{\prime},-\omega_{S})\mathcal{E}_{j}(\omega_{L})\mathcal{E}_{k}(\omega_{L}^{\prime})E_{l}^{\dagger}(\omega_{S}),$ (1)
[…]
$\Psi_{li}(\omega_{S},\omega_{aS})=C\mathcal{E}_{0j}\mathcal{E}_{0k}\int_{-\omega_{S}}^{\infty}G(\omega_{S}+\omega)G(\omega_{aS}-\omega)$
×$\times\left[A^{E}_{ijkl}+A^{R}_{ijkl}\frac{\gamma}{\omega_{ph}-\omega+i\gamma/2}\right]d\omega,$ (2)
where the total SaS susceptibility (in brackets) is composed of a constant electronic and a resonant phononic component [5, 14, 4].
The diamond structure belongs to the $O_{h}$ point group [13, 14]. In this case, $\chi^{(3)}_{ijkl}$ has only four independent terms [6], which are $\chi^{(3)}_{xxxx},\chi^{(3)}_{xxyy},\chi^{(3)}_{xyxy},\chi^{(3)}_{xyyx}$ (assuming light propagating in the $z$ direction and that the $x$ and $y$ directions coincide with the crystallographic axes). The frequency dispersion of the electronic contribution e-SaS can be neglected when it is very far from any electronic resonance, so that $\chi^{(3)E}_{ijkl}(-\omega_{aS},\omega_{L},\omega_{L}^{\prime},-\omega_{S})=A^{E}_{ijkl}\delta(\omega_{L}+\omega_{L}^{\prime}-\omega_{S}-\omega_{aS})$ becomes a complex constant tensor fulfilling energy conservation. Ideally, the ratio between the e-SaS components in centrosymmetric materials is $\chi^{(3)E}_{xxxx}\approx 3\chi^{(3)E}_{xyyx}\approx 3\chi^{(3)E}_{xxyy}\approx 3\chi^{(3)E}_{xyxy}$ [4]. However, experimental conditions are rarely ideal and we will only take these values as a reference.
The p-SaS susceptibility tensor, on the other hand, is known to be $\chi^{(3)R}_{ijkl}\propto\sum_{\sigma}(\alpha^{R}_{ij,\sigma}\alpha^{R}_{kl,\sigma}+\alpha^{R}_{ik,\sigma}\alpha^{R}_{jl,\sigma})$ [13], where $\alpha^{R}_{ij,\sigma}$ is the polarizability Raman tensor that describes the scattering of an incident electric field polarized at $i$ into a scattered mode polarized at $j$, via a phonon $\sigma$. The Raman active vibrational mode of diamond belongs to the $T_{2g}$ irreducible representation of the $O_{h}$ point group. Due to the relatively small probability of formation of SaS pairs, the spectral features of the phononic contribution to the susceptibility can be calculated from perturbation theory [4], and $A^{R}_{ijkl}$ in Eq. (2) is the tensorial part of $\chi^{(3)R}_{ijkl}$. The frequency dependence is composed of a Lorentzian-shaped amplitude of the Stokes scattering phonon frequency, $\omega=\omega_{L}-\omega_{S}$, around the resonance $\omega_{ph}$ with width $\gamma/2$ ($\gamma$ is the phonon decay rate, inversely proportional to its lifetime). This susceptibility is the same as when the scattered fields are classical [5, 4], being an extrapolation of the stimulated Raman scattering to the regime where the Stokes stimulation is turned off.
The SaS scattering amplitude is widened by the laser bandwidth, which adds an energy uncertainty to the scattered photons in both electronic and phononic processes. The factor $C$ in Eq. (2) contains the efficiency of the SaS scattering in polarizations $l$ (Stokes) and $i$ (anti-Stokes), and depends on the scattered field frequencies and the scattering angles. However, we can consider $C$ a constant in our experiments because we only collect pairs in forward scattering and because the laser bandwidth and phonon spectrum, taken care separately in Eq. (2), dominate any other frequency dependence. Assuming a Gaussian amplitude spectrum for the laser, $G(\omega_{L})=(\pi W^{2})^{-1/4}e^{-(\omega_{L}-\omega_{c})^{2}/2W^{2}}$, centered around the angular frequency $\omega_{c}$ with width $W$ (the laser power spectrum FWHM is $2\sqrt{\text{ln}2}W/(2\pi)$), the integral (2) can be solved analytically.
With the probability amplitude (2) we can construct the SaS two-photon state generated by the scattering process. We assume the laser polarization to be vertical (V) with respect to the laboratory, denoted by $\hat{\mathbf{e}}_{V}$. If the crystallographic axis $x$ is also along the vertical direction, then $\hat{\mathbf{e}}_{V}=\hat{\mathbf{x}}$, and the susceptibility $\chi^{(3)}_{ixxl}$ will govern the photon pair production. In general, however, if the crystallographic axis $x$ and the laser polarization have an angle $\theta$ between them, such that $\hat{\mathbf{e}}_{V}=\cos(\theta)\hat{\mathbf{x}}+\sin(\theta)\hat{\mathbf{y}}$, there will be a coherent sum of certain components of the susceptibility tensor defining the scattered photons polarization. By writing the laser field components as $\mathcal{E}_{0x}=\mathcal{E}_{0}\cos\theta$ and $\mathcal{E}_{0y}=\mathcal{E}_{0}\sin\theta$, the scattered (non-normalized) state can be written as
$|\Psi_{SaS}\rangle(\omega_{S},\omega_{aS};\theta)=[Y^{E}_{VV}(\theta)f^{E}+Y^{R}_{VV}(\theta)f^{R}]|VV\rangle$
$+[Y^{E}_{HH}(\theta)f^{E}+Y^{R}_{HH}(\theta)f^{R}]|HH\rangle$
$+[Y^{E}_{VH}(\theta)f^{E}+Y^{R}_{VH}(\theta)f^{R}](|VH\rangle+|HV\rangle),$ (3)
where $f^{\eta}(\omega_{S},\omega_{aS})$ carries the spectral dependence of the scattering for $\eta=\{E,R\}$, being $f^{E}\equiv e^{-\bar{\omega}^{2}/W^{2}}$ and
$f^{R}\equiv e^{-\bar{\omega}^{2}/W^{2}}\frac{\gamma e^{-(\Omega-i\gamma/2)^{2}/W^{2}}}{2i\sqrt{\pi}W}\text{erfc}\left(\frac{\gamma}{2W}+i\frac{\Omega}{W}\right),$ (4)
with $\bar{\omega}\equiv(\omega_{S}+\omega_{aS})/2-\omega_{c}$ and $\Omega\equiv(\omega_{aS}-\omega_{S})/2-\omega_{ph}$. Note that Eq. (4) comes from the integral in Eq. (2), and the complex denominator in the Raman term yields the complex argument in the error function, associated with the phonon decay. Also note that in $f^{R}$ the $e^{-\bar{\omega}^{2}/W^{2}}$ factor implied that detections at $\omega_{S}+\omega_{aS}=2\omega_{c}$ are favored, while the dependence with $\Omega$ means that what dictates the spectrum is the SaS combined frequency $(\omega_{aS}-\omega_{S})/2$ around $\omega_{ph}$. The functions $Y^{\eta}_{li}(\theta)$ depend on the angle $\theta$ and the $A^{\eta}_{ijkl}$ tensor components,
$Y^{\eta}_{VV}(\theta)$$\equiv$ $C\mathcal{E}_{0}^{2}[(\mathcal{S}^{4}+\mathcal{C}^{4})A^{\eta}_{xxxx}$ (5a)
$+2\mathcal{S}^{2}\mathcal{C}^{2}(A^{\eta}_{xyyx}+A^{\eta}_{xxyy}+A^{\eta}_{xyxy})],$
$Y^{\eta}_{HH}(\theta)$$\equiv$ $C\mathcal{E}_{0}^{2}[2\mathcal{S}^{2}\mathcal{C}^{2}A^{\eta}_{xxxx}+(\mathcal{S}^{4}+\mathcal{C}^{4})A^{\eta}_{xyyx}$ (5b)
$-2\mathcal{S}^{2}\mathcal{C}^{2}(A^{\eta}_{xxyy}+A^{\eta}_{xyxy})],$
$Y^{\eta}_{VH}(\theta)$$\equiv$ $C\mathcal{E}_{0}^{2}(\mathcal{S}^{2}-\mathcal{C}^{2})\mathcal{S}\mathcal{C}$ (5c)
×$\times\left(A^{\eta}_{xxxx}-A^{\eta}_{xyyx}-A^{\eta}_{xxyy}-A^{\eta}_{xyxy}\right),$
where $\mathcal{S}\equiv\sin\theta$ and $\mathcal{C}\equiv\cos\theta$. Equation (3) means that the angle $\theta$ drives the balance between the contribution of the e-SaS spectrum in $f^{E}$ and the p-SaS spectrum in $f^{R}$, and this balance will be different for each pair of scattered photons polarization $VV$, $HH$ and $VH$ (equal to $HV$), depending on how the tensor components are combined. If a different, possibly more complicated, laser spectrum were used, a new spectral function $f^{R}$ would result, but the state of Eq. (3) would still have the same form.
The state in Eq. (3) is, in general, entangled in polarization. In particular, for $\theta=0^{\circ}$, $Y^{\eta}_{VH}(0^{\circ})=0$, and $\{|VV\rangle,|HH\rangle\}$ form a Schmidt basis [17], meaning that if the amplitudes of these vector components are equal, state $|\Psi_{SaS}\rangle$ (3) is maximally entangled, while if one of them is zero, it is separable. In this sense, the problem of producing a state with a high amount of polarization entanglement becomes a matter of tailoring the relative amplitude of the two components of this Schmidt basis.
To obtain a prediction for the SaS state generated as a function of $\theta$, one needs to characterize the tensor components $A^{\eta}_{ijkl}$ in the functions $Y^{\eta}_{li}(\theta)$, Eq. (5). The functions $f^{\eta}$ in Eq. (3) are known and only depend on the laser central frequency $\omega_{c}$, bandwidth $W$, the phonon frequency $\omega_{ph}$ and decay rate $\gamma$, which are given experimental conditions. Measuring the relative scattered intensities into $|HH\rangle$ and $|VV\rangle$ in $\theta=0^{\circ}$ and $\theta=45^{\circ}$ (angles in which there are no $|VH\rangle$ and $|HV\rangle$ components) allows us to retrieve the values of $A^{\eta}_{ijkl}$. Next, we present an experiment that was used to extract these quantities in diamond, and show how they translate into SaS entanglement.
[…]
The experimental setup utilized is similar to the one described in [12], with the addition of a monochromator to filter the Stokes signal [18]. The polarization of the laser is vertical (V) with respect to the laboratory and, for each monochromator position (each selected $\delta\omega=\omega_{c}-\omega_{S}$), one histogram of temporal difference is obtained in $300$ seconds of acquisition. This scanning procedure is done for two orientations of the sample ($\theta=0^{\circ}$ and $45^{\circ}$), and for each sample orientation the spectrum is obtained with the polarization of the photons incident on the avalanche photodiode (APD) detectors being selected in two ways—a spectrum with Stokes and anti-Stokes photons with vertical polarization (VV) and a spectrum with both photons with horizontal polarization (HH). The sample utilized was a diamond grown by a CVD process (Type IIac, 100-oriented, from Almax) positioned so that the laser propagates in the (001) direction of the crystal.
The intensity spectra of correlated SaS pairs for VV (blue dots) and HH (red dots) polarization at $\theta=0^{\circ}$ and $\theta=45^{\circ}$ are displayed in Fig. 1 (a) and (b), respectively. This experiment was specifically designed by combining the single-wavelength polarization analysis presented in [12] with the unpolarized spectral analysis provided in [15], to prove the strength of our theoretical approach. According to group theory analysis, VV(0) configuration does not exhibit a Raman contribution ($A^{R}_{xxxx}=0$ in Eq. (5)) and only involves non-resonant electronic transitions. For this reason, the blue curve shown in Fig. 1 (a) corresponding to this configuration is flat. Additionally, Fig. 1 (a) shows a light blue solid line, obtained by averaging the intensity values $I_{SaS}^{corr}$ for this experimental configuration, $\langle I_{SaS}(VV(0^{\circ})\rangle=(27.5\pm 3.5)\times 10^{3}$. Since $Y^{E}_{VV}(0^{\circ})=C\mathcal{E}_{0}^{2}A^{E}_{xxxx}$ and in our model it is proportional to $\langle I_{SaS}(VV(0^{\circ})\rangle^{1/2}$, we have a reference value for $A^{E}_{xxxx}$.
By looking at the other relevant scattering geometries, we obtain the other tensorial components of Eq. (5) in relation to $A^{E}_{xxxx}$. The HH(45) configuration also exhibits only e-SaS, as seen in the red data in Fig 1 (b). The mean and standard deviation of the counts gives $(1.60\pm 0.28)\times 10^{3}$. The HH(0) data in Fig. 1 (a) (red points) contains p-SaS with the characteristic resonance signature, which comes from $A^{R}_{xyyx}\neq 0$ in Eq. (5b), contributing with a Raman scattering in the $|HH\rangle$ polarization. Using $Y^{E}_{HH}(0^{\circ})=\langle I_{SaS}^{corr}(\text{VV}(0^{\circ}))\rangle^{1/2}(0.68-i0.12)$ and $Y^{R}_{HH}(0^{\circ})=51450$ fits the data with the shown theoretical curve in Fig. 1 (a). The VV(45) data is the last configuration left, and it is shown in Fig. 1 (b) in blue. In order to fit it, we fix the Raman factor to be the same as in the HH(0) configuration, $Y^{R}_{VV}(45^{\circ})=51450$, and use $Y^{E}_{VV}(45^{\circ})=\langle I_{SaS}^{corr}(\text{VV}(0^{\circ}))\rangle^{1/2}(1.61-i0.55)$. With the above values and working with Eq. (5) we obtain the electronic and Raman $A^{\eta}_{ijkl}$ components, summarized in Table 1.
Table 1: List of measured electronic and Raman tensor components ratios $A^{\eta}_{ijkl}/A^{E}_{xxxx}$ [18].
Tensor component Value
$A^{E}_{xyyx}/A^{E}_{xxxx}$ $(0.37\pm 0.04)+i(-0.07\pm 0.01)$
$(A^{E}_{xxyy}+A^{E}_{xyxy})/A^{E}_{xxxx}$ $(0.89\pm 0.0.10)+i(-0.07\pm 0.01)$
$A^{R}_{xxxx}/A^{E}_{xxxx}$ 00
$A^{R}_{xyyx}/A^{E}_{xxxx}$ $(171\pm 12)$
$(A^{R}_{xxyy}+A^{R}_{xyxy})/A^{E}_{xxxx}$ $(171\pm 12)$
For completeness the second-order correlation function $g^{(2)}(0)$ is evaluated with the ratio $[I^{corr}_{SaS}(\Delta\tau=0)+\bar{I}_{SaS}(\Delta\tau\neq 0)]/\bar{I}_{SaS}(\Delta\tau\neq 0)$, where $I^{corr}_{SaS}(\Delta\tau=0)$ are the measurements in Fig. 1 (a) and (b), and $\bar{I}_{SaS}(\Delta\tau\neq 0)$ accounts for the uncorrelated pair production. A plot of $g^{(2)}(0)$ is shown in Fig. 1(c,d), where it is evident that near the resonance the values are the lowest, accounting for the high number of uncorrelated SaS photon pairs produced. The large error bars in $g^{(2)}(0)$ reflect the errors in the low accidental counts (see [18]), and the theory agrees well with the data in spite of the error size. Furthermore, the curves for HH(0) and VV(45) show a drop in $g^{(2)}(0)$, being higher below and lower above the resonance region, while the VV($0^{\circ}$) and HH($45^{\circ}$) keep roughly the same value. The asymmetry in HH(0) and VV(45) is associated with the uncorrelated counts being symmetrical with respect to the resonance peak, and thus the asymmetry in the correlated counts for these configurations, which contain a p-SaS contribution, is transferred to the $g^{(2)}(0)$ curves. This asymmetry is a result of the constructive ($\delta\omega<\omega_{ph}$) versus destructive ($\delta\omega>\omega_{ph}$) interference between the e-SaS and the p-SaS, which explains the Cooper-pair-like behavior of $I_{SaS}$ [10, 15]. On the other hand, the correlated counts in the VV(0) and HH(45) configurations are a purely e-SaS contribution, and thus are a flat curve, leading to an overall symmetric $g^{(2)}(0)$.
With the values of $A^{\eta}_{ijkl}$ in hand, one can predict what is the entanglement in the SaS scattered state $|\Psi_{SaS}\rangle(\theta)$ of Eq. (3) for given values of $\omega_{S}$ and $\omega_{aS}$. We use, as an entanglement measure, the entropy of entanglement $E=-\text{Tr}_{i}(\rho_{i}\text{log}_{2}\rho_{i})$ of either subsystem $i=\{S,aS\}$ with reduced state $\rho_{i}$ [17], which, since our global state is pure, is zero for separable states and unity for maximally entangled states.
Figure 2: Entanglement map for the SaS states in (a) the present experiment and (b) a generic centrosimetric cubic system. In (a), the laser bandwidth $W$ is fixed at $W=264\text{ cm}^{-1}$, the value of our experiment, as localized in (b) by a horizontal line. The black vertical line is the resonant Raman shift $\omega_{ph}/2\pi$, and the red regions mark the entanglement maxima ($E>0.999$). As tensor components in Eq. (5) we used the values from table 1. The star localizes the experiment of Ref. [12]. In blue shades, we plot contour curves for the Gisin parameter $F$. The hatched regions, within one laser bandwidth FWHM, indicates a high production of independent SaS photon pairs.
In Fig. 2 we show a contour plot of $E(\omega_{S},\omega_{aS};\theta,W)$, under the condition that $\omega_{aS}=2\omega_{c}-\omega_{S}$, that is, a symmetric SaS Raman shift. On the horizontal axis we vary the Raman shift, going through the resonance in diamond at $\omega_{ph}/2\pi=1332$ cm-1, identified by a black vertical line, and on the vertical axis we plot (a) the angle $\theta$ between the laser linear polarization and the crystallographic axis, going from $\theta=0$ to $45^{\circ}$ (due to the symmetry of the crystal, all other $\theta$ values will be related to this range), and (b) the laser bandwidth $W$, going from zero to $W=190\gamma$, and where the value for our experiment $W=24\gamma$ ($W=264\text{ cm}^{-1}$ or FWHM 70 cm-1) is indicated by a black horizontal line. The thick red regions indicate the parameters where $E\rightarrow 1$, showing maximum entanglement.
Maximum entanglement occurs at $\theta=0^{\circ}$, in a region below resonance close to 1200 cm-1 and another above, close to 1400 cm-1. The measurements in Fig. 1 illustrate this for the $0^{\circ}$ and $45^{\circ}$ conditions, for which $|HH\rangle$ and $|VV\rangle$ are a Schmidt basis, and the balance between $HH$ and $VV$ counts reflects how close it is to maximum entanglement. At $0^{\circ}$ the purely electronic response (VV) is strong, and maximum entanglement occurs where the HH response crosses it with the same amount of scattered photon pairs. Conversely, at 45 the purely electronic response (HH) is weak in comparison with the VV scattering, which contains the Raman response, thus making it impossible for the probability amplitude of the $|HH\rangle$ component to balance with the $|VV\rangle$ one, when it would reach a maximally entangled state. For other crystal angles $\theta$, comprising the middle section of Fig. 1 (a), the Schmidt basis can be determined for any $\theta$ using our model of Eq. (3). Measurements on this basis allows one to experimentally obtain the SaS pair entanglement entropy for any desired crystal angle without the need for a complete state tomography.
If the laser bandwidth $W$ grows, as shown in Fig. 2 (b), the shape of the resonance curve is smoothed out, so the ratio between the $|VV\rangle$ and $|HH\rangle$ coefficients gets more and more even along the SaS spectrum. Because of this, the entanglement extremes become less pronounced as $W$ grows, until the point that it is not possible to reach maximum entanglement anymore, at $W\approx 70\gamma$. On the other hand, it also gets harder to obtain a separable state, which depends on a peak of either the $|VV\rangle$ or the $|HH\rangle$ component in relation to the other.
In Fig. 2 there is a hatched region indicating where the uncorrelated SaS pair production is high (low $g^{(2)}$ region in Fig. 1 (c,d)), corresponding to $2W$, around 1.2×1.2$\times$1.2 × FWHM. In this region, the scattered state is not properly represented by Eq. (2), but it needs to be complemented by other SaS scattering events, which involves the scattering of real phonons, that happens within a laser bandwidth around the Raman resonance [19]. Outside this region, though, state (2) is a good representation of the scattered SaS state, and our entanglement measure is representative.
To complete the analysis, we have also drawn level curves for the violation of Bell-type CHSH inequalities as a function of the light-matter parameters. These curves correspond to the so-called Gisin parameter $F$ [20, 21] which, for pure states, reads $F=2(1-\mathcal{P})^{1/2}$, where $\mathcal{P}=1-\text{Tr}_{i}(\rho_{i})^{2}$ is the linear entropy of either subsystem $i$. For separable states $F=2$, which means that no separable state can go above the classical upper bound in the CHSH inequality, equal to 2. It reaches its maximum at $F=2\sqrt{2}\approx 2.83$ for maximally entangled states. We plot the $F$ parameter in Fig. 2 as blue shaded contour curves, and its smallest values within a white contour. Note that the maximal violation coincides with the maximally entangled states near 1200 cm-1 and 1400 cm-1. Since Bell analysis is state dependent, and the generated $|\Psi_{SaS}\rangle$ state depends on the values of the light-matter parameters, the appropriate Bell angles to reach the maximum violation change for different regions of the map. As an example, the violation in the maximally entangled states near 1200 cm-1 is obtained for linear polarization angles $(0,\pi/4)$ for one of the photons and $(\pi/8,3\pi/8)$ for the other, where the angles are in relation to $V$ polarization. Note that the quality of the entanglement, characterized by both the Gisin parameter $F$ and the entanglement entropy $E$, are independent of the rate of pair generation, so no matter how many SaS photon pairs are generated, our analysis yields how much entanglement will be available with each pair.
Finally, we can localize the result of Ref. [12] in Fig. 2, represented by a white star. In the paper, it was found a state close to a Bell state at a symmetric Raman shift of 900 cm-1 with $\theta=0^{\circ}$, with amplitude ratio of $|HH\rangle$ to $|VV\rangle$ at $\sqrt{0.28/0.72}\approx 0.62$. Calculating the ratio with our state (3) yields 0.49, and the discrepancy comes from our aS filter cutting some of the Raman contribution in HH at 900 cm-1 (left edge of Fig. 1 (a)), which does not happen in Ref. [12]. This region can be seen in the bottom left of Fig. 2 (a) and (b), where and $E\approx 0.7$ and $F\approx 2.5$.
To conclude, this work presents a fully quantum theory that describes the broadband scattered state in the SaS process, including photon pair polarization. It incorporates experimental conditions like the laser bandwidth and crystal orientation, to allow for the calculation of any sought physical quantity (e.g. $g^{(2)}(0)$ and entanglement). The efficiency of the process can be controlled by how close the system is to the resonance with a Raman active phonon, and it is a matter of state engineering to choose the appropriate laser–crystal angle and SaS frequency in order to obtain a good balance between a good amount of entanglement and a sufficient pair production rate. This theory has been elusive to find despite a decade-long effort [2, 10, 15, 22, 23, 12]. The result is explored here for diamond, but it should be general for other centrosymmetric media, including silicon, whose electronic band gap (1.12 eV) is one order of magnitude higher than the phonon energy ($<0.1$ eV), respecting the assumptions of our model. This should considerably bolster the engineering of SaS photon pair generation.
Other good candidates for SaS photon pair generation which can feed on the principles of our theory are two-dimensional materials. Graphene, however, is limited by a strong luminescence that prevents photon correlation studies [24, 2], and in transition metal dichalcogenides (TMDs) the phonon energies are too low, which demands low cryogenic temperatures to avoid a high rate of thermally excited anti-Stokes photons.
For nonzero $\chi^{(2)}$ materials, spontaneous parametric down-conversion (SPDC) enables broadband pair generation as well, as it has been recently demonstrated [25]. However, in SPDC the generation of a polarization entangled photon pair is due to the decay of a single highly energetic photon [5], demanding a UV laser for the generation of photon pairs in the visible spectrum, while the use of FWM allows the generation of pairs with frequencies close to any excitation laser. This peculiarity means that using FWM can have more versatility in the choice of optics for entangled photon pair generation.
Acknowledgments
This work was supported by IDOR/Pioneer Science Initiative (www.pioneerscience.org), CNPq (INCT-IQ 465469/2014-0, 302872/2019-1, 421469/2023-4, 307619/2023-0), FAPEMIG (APQ-01860-22, APQ-04852-23, RED0008123) and FAPERJ (CNE E-26/200.307/2023).